### v3

# InfiniteHBD: Building Datacenter-Scale High-Bandwidth Domain for LLM with Optical Circuit Switching Transceivers

Abstract.
Scaling Large Language Model (LLM) training relies on multi-dimensional parallelism, where High-Bandwidth Domains (HBDs) are critical for communication-intensive parallelism like Tensor Parallelism (TP) and Expert Parallelism (EP). However, existing HBD architectures face fundamental limitations in scalability, cost, and fault resiliency: switch-centric HBDs (e.g., NVL-72) incur prohibitive scaling costs, while GPU-centric HBDs (e.g., TPUv3/Dojo) suffer from severe fault propagation. Switch-GPU hybrid HBDs such as TPUv4 takes a middle-ground approach by leveraging Optical Circuit Switches, but the fault explosion radius remains large at the cube level (e.g., 64 TPUs).
We propose InfiniteHBD, a novel transceiver-centric HBD architecture that unifies connectivity and dynamic switching at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, InfiniteHBD achieves reconfigurable point-to-multipoint connectivity, allowing the topology to adapt into variable-size rings. This design provides: i) datacenter-wide scalability without cost explosion; ii) fault resilience by isolating failures to a single node, and iii) full bandwidth utilization for fault-free GPUs. Key innovations include a Silicon Photonic (SiPh) based low-cost OCS transceiver (OCSTrx), a reconfigurable k-hop ring topology co-designed with intra-/inter-node communication, and an HBD-DCN orchestration algorithm maximizing GPU utilization while minimizing cross-ToR datacenter network traffic. The evaluation demonstrates that InfiniteHBD achieves 31% of the cost of NVL-72, near-zero GPU waste ratio (over one order of magnitude lower than NVL-72 and TPUv4), near-zero cross-ToR traffic when node fault ratios under 7%, and improves Model FLOPs Utilization by 3.37x compared to NVIDIA DGX (8 GPUs per Node).
copyright: nonedoi: isbn: conference: ; 2025journalyear: ;price: 11footnotetext: Guyue Liu, Hao Nie, and Huaiyu Meng contributed equally to this work and share the corresponding authorship.
1. Introduction
Large-scale Large Language Models (LLMs) training rely on various parallelism strategies (Smith et al., 2022; Rajbhandari et al., 2020), such as Tensor Parallelism (TP), Expert Parallelism (EP), Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP) and Sequence Parallelism (SP). These strategies communicate over two types of AI datacenter compute fabrics, each with distinct bandwidth requirements. First, Datacenter Networks (DCNs) provide hundreds of Gbps per GPU and primarily handle DP, PP, CP, and SP traffic, which has lower communication demands. Second, High-Bandwidth Domains (HBDs) offer Tbps-level bandwidth, which is crucial for communication-intensive TP and EP. Efficient HBD design can reduce communication overhead, thereby improving Model FLOPs Utilization (MFU)-a key performance metric for LLM training.
The community has made significant advancements in designing DCNs for LLM training (Wang et al., 2024; Qian et al., 2024; NVIDIA, 2021; Gangidi et al., 2024a). However, scaling HBD to optimize MFU in LLM training remains a challenging problem. Existing HBD architectures (NVIDIA, 2024b; Jouppi et al., 2023; Tesla, 2024; Jouppi et al., 2020; Trainium, 2024) take important steps but still suffer from fundamental limitations in scalability, cost, and fault resiliency.
  • $\bullet$
    Switch-centric HBDs, such as NVIDIA NVL-72 (NVIDIA, 2024b), build multilayer nonconvergent networks for HBD with switch chips. However, the switch fabric incurs superlinear cost growth as it scales, constraining the number of GPUs per HBD. This limitation prevents optimal large TP and EP and causes severe resource fragmentation when the size of TP/EP group increases. For instance, with 2 HBDs (32 GPUs each), 30 GPUs are wasted for TP-16 jobs if each HBD has a single GPU failure. This waste reduces to 14 GPUs if the two HBDs are combined into a 64-GPU unit.
  • $\bullet$
    GPU-centric HBDs, such as Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021), adopts low-cost GPU-to-GPU links to construct large-scale ring or mesh topologies, forwarding traffic directly through GPUs. However, these architectures suffer from a large fault explosion radius, where a single GPU failure degrades bandwidth for a group of adjacent GPUs, compromising the entire topology. For example, in SiP-Ring, one single GPU failure breaks the ring and reforms the topology a line.
  • $\bullet$
    Switch-GPU Hybrid HBDs, TPUv4 (Jouppi et al., 2023) alleviates the limitation of large fault explosion radius via OCS-based switch [^1] In this paper, ”OCS” specifically denotes optical circuit switching capability, while OCS-based switch denotes optical circuit switch.: each set of 64 TPUs is connected as a cube, with these cubes connected to multiple OCS-based switches to isolate faults within the cubes. However, it cannot resolve it fundamentally, as the fault explosion radius remains large at the cube level (64 TPUs).
In this paper, we take a first-principles approach to redesigning HBD for LLM training workloads. Through a top-down analysis of parallelism strategies (§2.3) for maximizing MFU, we show that increasing TP size yields the most significant MFU gains for both dense and sparse LLM models. For large dense models (Llama Team, AI @ Meta, 2024), the optimal TP size scales from 16 to 64 as the number of GPUs increases. For sparse MoE models (Sun et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024b), enlarging TP improves MFU more effectively than EP, particularly when considering the expert imbalance problem (Liu et al., 2023b).
These findings lead to two key design principles: i) HBD should be optimized exclusively for TP Ring-Allreduce communication with large message sizes, which communicates with only logical neighboring nodes, eliminating the need for EP and the associated any-to-any communication; ii) Supporting large and adaptable TP is essential, as different GPU numbers and model sizes require varying TP configurations to maximize MFU.
Based on these principles, we propose InfiniteHBD, a scalable and fault-resilient HBD architecture designed for optimizing TP communication. Our key insight is unifying connectivity and dynamic switching at the transceiver level using OCS. By embedding OCS in each transceiver, we achieve reconfigurable point-to-multipoint connectivity. This marks a departure from traditional designs, where transceivers support only point-to-point connections and rely on high-radix switches for routing. We call this new design transceiver-centric HBD architecture. This transceiver-centric architecture offers two key benefits: i) It enables the flexible construction of arbitrarily large ring topologies by intra-node loopback mechanism. This can support optimal TP group sizes for different models, while effectively minimizing resource fragmentation; ii) When one node fails, its neighboring transceivers dynamically reconfigure connections to reroute traffic, significantly reducing the fault explosion radius and improving system resilience.
We realize the transceiver-centric HBD architecture in production by combining the following key ideas:
  • $\bullet$
    Silicon Photonics based OCS transceiver (OCSTrx): To design a cost-effective low-power transceiver with OCS support, we leverage the current advances of Silicon Photonics (SiPh) technology. Compared to MEMS (Urata et al., 2022a; Yeow et al., 2001) technology which has been widly used to realize OCS, SiPh offers simpler structures, lower cost and power consumption. We build OCS with Mach-Zehnder interferometer (MZI) matrix (Wikipedia, 2024), taped out with $65nm$ CMOS processes. The chip size is less than $136.5mm^{2}$ while the chip power consumption is $3.2Watts$, which can be integrated into commercial QSFP-DD $800Gbps$ transceiver (MSA, 2024) with sub-1ms path reconfiguration latency.
  • $\bullet$
    Reconfigurable K-Hop Ring Topology: While OCSTrx offers reconfigurable connections at the transceiver level, constructing adaptive-size rings that maximize GPU utilization remains a challenge. For example, a naive full-mesh topology built with OCSTrx would impose strict limits on TP size, while also resulting in significant bandwidth waste and fragmentation. To address this, we propose a reconfigurable K-Hop Ring, where each node connects to all other nodes within $\leq K$ hops via OCSTrx. The intra-node loopback mechanism enables dynamic ring construction, while the inter-node backup link bypasses faulty nodes, ensuring high fault tolerance.
  • $\bullet$
    HBD-DCN Orchestration Algorithm: While an optimal HBD topology is critical, end-to-end training performance also depends on efficient HBD-DCN coordination. For example, the orchestration of TP groups in HBD directly determines DP traffic distribution, which impacts congestion in DCN, ultimately governing training performance. Unfortunately, existing approaches lack mechanisms to jointly coordinate DCN and HBD to mitigate congestion and optimize communication efficiency. To address this, we propose a new orchestration algorithm that minimizes cross-ToR traffic, thereby minimizing congested traffic.
To the best of our knowledge, InfiniteHBD is the first HBD design capable of scaling to datacenter scale while avoiding cost explosion and increased failure-induced waste. We evaluated InfiniteHBD with the real 348-day fault trace from our 10K GPU cluster [^2] Details in Appendix §A. We have open-sourced this trace at https://gitlab.basemind.com/sys/infra/superpod/infinitehbd-trace. When executing TP32 jobs with the trace, it demonstrates 0.15% GPU waste ratio - 69x and 45x lower than NVL-72 (10.31%) and TPUv4 (6.76%). It achieves 3.24x and 1.59x cost reductions compared to NVIDIA NVL-72 and Google TPUv4 respectively. Through the orchestration algorithm, it maintains near-zero cross-ToR traffic under 7% node failure rates. Its dynamic ring formation capability enables 3.37x higher MFU than NVIDIA DGX systems (Platform, 2024) (8 GPUs/node).
2. Background and Motivation
In this section, we first introduces LLM training in AI datacenters (DCs) (§2.1). Then, we examine existing High-Bandwidth Domain (HBD) architectures and discuss their limitations (§2.2). Finally, we summarize key design principles of HBD for LLM training (§2.3).
2.1. LLM Training in AI DC
(a) Switch-centric: NVL36
(b) GPU-centric: SiP-Ring
(c) GPU-centric: Dojo
(d) Hybrid: TPUv4
Figure 1. Illustrative examples of HBD architectures. N represents Node, and S represents Switch. Red (with cross hatch) represents fault device and yellow (with dots) represents unavailable or downgraded GPU.
Architecture Type Scalability Collective Primitives Fault Explosion Radius Interconnect Cost Fragmentation
Node-Side Switch-Side
NVL Switch-centric Low Full CCL Node-level Switch-level High Many
Dojo, TPUv3, SiP-Ring GPU-centric High Ring-Allreduce HBD-level Low Few
TPUv4, TPUv5p Switch-GPU Hybrid Moderate Ring-Allreduce Cube-level Switch-level Moderate Few
InfiniteHBD Transceiver-centric High Ring-Allreduce Node-level Low Few
Table 1. Comparative analysis of HBD architectures.
LLM training parallelism and communication. LLM training jobs employ various parallelism strategies to efficiently utilize GPUs distributed across AI DCs (Shoeybi et al., 2020; Rajbhandari et al., 2020). Based on communication loads, parallelism can be categorized into two types. The first type is communication-intensive parallelism which involves high communication load. Tensor Parallelism (TP) splits the model across multiple GPUs and synchronizes via AllReduce. The ring algorithm for AllReduce is theoretically optimal (Patarasuk and Yuan, 2009), making ring-based topologies ideal for TP. Expert Parallelism (EP), designed for Mixture of Experts (MoE) models (Sun et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024b), assigns experts to different GPUs and relies on AlltoAll communication, requiring topologies with high bisection bandwidth (e.g., Full-Mesh). In contrast, parallelism strategies such as Data Parallelism (DP), Pipeline Parallelism (PP), Context Parallelism (CP), and Sequence Parallelism (SP) introduce lower communication overhead, placing less demands on network performance.
Compute fabric. Compute fabric in AI DC interconnects GPUs to efficiently transmit model gradients and parameters. It consists of two primary components: Datacenter Network (DCN) and High-Bandwidth Domain (HBD). DCN provides communication across the entire AI DC via Ethernet or Infiniband, the bandwidth is around $200\sim 800Gbps$. Widely used DCN architectures include Fat-Tree (Al-Fares et al., 2008) and Rail-Optimized (NVIDIA, 2021). In comparison, HBD offers Tbps-level throughput, and is more suitable for TP/EP. However, its scale is typically constrained by interconnection costs and fault tolerance considerations. For example, NVL-72 (NVIDIA, 2024b) only interconnects 72 GPUs per HBD.
Faults and fault explosion radius. As revealed by current advances of AI DCs (Qian et al., 2024; Gangidi et al., 2024b), training jobs experience a variety of faults, such as GPU faults, optical transceiver faults, switch faults, and link faults. We quantify the fault impact using the fault explosion radius, defined as the number of GPUs degraded by a single fault event. The fault explosion radius varies depending on both the system architecture and the fault component. For example, if a switch fails, the bandwidth of all devices connected to it will degrade, illustrating the switch-level fault explosion radius.
HBD fragmentation. When the number of GPUs in the HBD cannot be evenly divided by the size of the parallel group (i.e., TP size), the remaining GPUs become unusable, leading to resource waste. The GPU waste ratio for each HBD can be expressed by the formula $\{(HBD_{size}-N_{fault})\mod TP_{size}\}/{HBD_{size}}$. In AI DCs with small-scale HBDs, GPU waste due to fragmentation is significant because each HBD experiences independent fragmentation. This issue worsens as the TP group size increases with model scale. For example, for NVL-36 shown in Figure 1(a), running TP-16 causes $\geq$11GPU waste ratio.
2.2. Limitations of Existing HBDs
Existing HBD architectures for LLM training can be categorized into three types, based on the key components that provide connectivity. A summary is shown in Table 1.
Switch-centric HBD. This type architecture leverages switch chips to interconnect GPUs, as shown in Figure 1(a). A prominent example is NVIDIA, which utilizes NVLink and NVLink Switch  (NVIDIA, 2024d, 2018), e.g. DGX H100 (Platform, 2024) with 8-GPU and GB200 NVL-36, NVL-72, and NVL-576 (NVIDIA, 2024b). These architectures offer high-performance any-to-any communication. However, switch-centric HBDs have several drawbacks: i) They require a large number of switch chips due to their limited per-chip throughput; ii) They are vulnerable to a switch-level fault explosion radius—when a switch chip fails, all connected nodes experience bandwidth degradation; iii) High interconnect costs constrain the scale of HBDs, leading to significant fragmentation when serving large models.
GPU-centric HBD. GPU-centric HBD architectures construct the HBD using direct GPU-to-GPU connections, eliminating the need for switch chips. As a result, cost scales linearly with HBD size. A representative example is SiP-Ring (Khani et al., 2021), shown in Figure 1(b), where GPUs are organized into fixed-size rings. However, this design imposes a strict limitation: the TP group size must remain fixed. To enable communication at dynamic scales and support a wider range of workloads, more complex topologies are adopted (e.g., Dojo (Tesla, 2024), NVIDIA V100 (Corporation, 2018), TPUv3 (Jouppi et al., 2020), and AWS Trainium  (Trainium, 2024) ), which support dynamic scaling by allowing jobs to execute on topology subsets of varying sizes. As shown in Figure 1(c), Dojo (Tesla, 2024) connects GPUs via mesh-like topologies and employ GPUs to forward traffic. While GPU-centric architectures mitigate cost explosion and can support various scales, they suffer from a large fault explosion radius. A single GPU failure can disrupt the entire HBD by altering its connectivity, degrading communication performance even for healthy GPUs—such as the yellow GPUs in Figure 1(c).
Switch-GPU Hybrid HBD. This architecture interconnects GPUs via a combination of direct GPU-to-GPU connections and switch links. A typical example is TPUv4 (Jouppi et al., 2023), which organizes TPUs into $4^{3}$ TPU cubes and connect them via centralized OCS-based switches (Figure 1(d)). TPUv4 scales up to 4,096 TPUs, with its expansion primarily limited by the port count of the OCS-based switch. Furthermore, it suffers from a cube-level fault explosion radius—a failure in any single TPU affects the entire 64-TPU cube, leading to significant performance degradation. Furthermore, OCS-based switches face challenges of high costs and manufacturing complexity, which undermines the cost-effectiveness of TPUv4. TPUv5p cluster (Google, 2024) is similar to TPUv4 but can scale out to 8,960 TPUs.
Figure 2. InfiniteHBD overview.
2.3. Key Attributes of An Ideal HBD
GPU TP PP DP MFU $\textbf{MFU}_{TP-8}$ Improve
1024 16 4 16 0.5236 0.5217 1.0036
4096 16 8 32 0.4668 0.4282 1.0901
8192 32 8 32 0.4247 0.3512 1.2093
16384 32 16 32 0.3756 0.2584 1.4536
32768 32 16 64 0.3090 0.1690 1.8284
65536 64 16 64 0.2493 0.0999 2.4955
131072 64 16 128 0.1851 0.0550 3.3655
Table 2. Optimal parallelism strategy for maximum MFU of Llama 3.1-405b, compared to the baseline MFU for TP-8 (e.g., in widely-deployed NVLink architectures), when GPU number varies.
Existing HBD architectures face fundamental limitations in interconnection cost, resource utilization, and failure resiliency when scaling. To guide a better design, we analyze existing training workloads and explore two key questions without the limitations imposed by current HBD: i) What is the optimal group size that HBD should support? ii) What traffic patterns should HBD accommodate?
Large and adaptable TP size is critical for dense models. The optimal LLM training parallelism depends on model architectures and cluster configurations. For example, as illustrated by previous work (Zhang et al., 2024; Zhong et al., 2024). We evaluate the Model FLOPs Utilization (MFU) for Llama 3.1-405B (most capable models to date, 2024) using our in-house LLM training simulator (§6.3) and report the results in Table 2. MFU and TP/PP/DP columns denote the optimal MFU when TP size is unconstrained and the corresponding parallelism strategies respectively. $MFU_{TP-8}$ column denotes the optimal MFU when TP size is limited to 8. As we increase the number of GPUs, the optimal TP size grows from 16 to 64, a trend we observe across other large dense models. In this case, the HBD scale restricts the maximum size of TP, which affects training performance as a result.
Parallelism Operation Traffic Load
TP AllReduce $2bsh\cdot\frac{n-1}{n}$
EP AllToAll $2bsh\cdot\frac{n-1}{n}\cdot\frac{k}{n}$
Table 3. Communication load of TP and EP on a single MoE layer. $b$: batch size; $s$: sequence length; $h$: hidden dim; $k$: topK of MoE router; $n$: parallel size. Assume each expert is assigned equal number of tokens.
MoE can also be efficient with large-size TP. Beyond widely used dense models, we also examine sparse MoE models, which are trending toward larger scales (e.g., 1T parameters (Fedus et al., 2022)). The distributed training for MoE can be achieved through TP or EP (or a combination of them) [^3] For TP, each expert is equally sharded to GPUs. For EP, each expert is indivisible and allocated to one GPU in the EP group. (Liu et al., 2023b), both TP and EP are communication-intensive (Li et al., 2023a), making them heavily reliant on HBD.
TP EP
imbalance coef - 0% 10% 20% 30%
MFU (%) 31.2 31.5 30.5 29.8 28.8
Table 4. Performance comparison of TP and EP when training GPT-MoE.
Our production training experience on a 1T MoE model in production brings the following insights into the pros and cons of TP and EP. On the one hand, EP is more communication-efficient than TP. Table 3 compares the communication volume of TP and EP. Clearly, EP is better if $k<n$, which is common (Liu et al., 2024) because existing models often choose small $k$ for higher computation sparsity. On the other hand, EP suffers from the well-known expert imbalance problem (Liu et al., 2023b), especially when the MoE routers use the no-token-left-behind algorithm (Liu et al., 2024; Gale et al., 2023; Du et al., 2022). This will result in non-equivalent number of tokens that each expert will receive, which hence causes straggler nodes that waste GPU cycles of other nodes. Table 4 shows the simulated result of training GPT-MoE with 1.1T parameters (details in Appendix §B) under different expert imbalance coefficients [^4] Calculated as $\frac{max-min}{max}$, where $max$ and $min$ represent the maximum and minimum tokens allocated to each expert respectively.. When $coef=0$, EP is better than TP due to smaller communication overhead. As $coef$ increases, the MFU drops because of the straggler issue.
Key findings. These experiments provide us two key findings for HBD design. First, larger HBD size is increasingly needed for rapidly scaling LLMs (i.e., more than 1T parameters). Second, with larger HBD enabled, using TP is more favorable than EP to train MoE, because TP shards the computation equally across GPUs and hence bypasses the expert imbalance problem.
These findings reveal two key design principles for HBD: i) HBD must inherently support large and adaptable TP sizes, which fundamentally requires the scalability of HBD architecture; ii) the HBD designs need to ensure the effective support for the Ring-AllReduce communication. Given the demonstrated efficiency of TP in MoE training, ensuring support for Ring-AllReduce support is sufficient for mainstream LLM training scenarios; iii) small fault explosion radius. Thus, we propose designing a large and adaptable HBD architecture tailored for ring-based TP communication to optimize LLM parallelism strategies.
3. Design Overview
In this section, we first present our new HBD architecture InfiniteHBD guided by the design principles outlined above. We then provide an overview of its key components.
(a) Components of OCS transceivers.
(b) Zoom into OCS MZI switch matrix.
Figure 3. Design of OCS Transceivers. The core component is OCS integrated in transceivers.
Transceiver-centric HBD architecture. As discussed in §2.2 and summarized in Table 1, existing architectures face a fundamental tradeoff among scalability, cost, and fault isolation. The GPU-centric architecture offers high scalability and low cost connectivity but suffers from a large fault explosion radius. In contrast, the switch-centric architecture improves fault isolation by leveraging centralized switches to confine failures to the node level. However, this comes at the cost of reduced scalability and higher connection overhead. The GPU-switch hybrid architecture takes a middle-ground approach but still suffers from significant fault propagation. As a result, no existing architecture fully meets all requirements.
Our key insight is that connectivity and dynamic switching can be unified at the transceiver level using Optical Circuit Switching (OCS). By embedding OCS within each transceiver, we can enable reconfigurable point-to-multipoint connectivity, effectively combining both connectivity and switching at the optical layer. This represents a fundamental departure from conventional designs, where transceivers are limited to static point-to-point links and rely on high-radix switches for dynamic switching. We refer to this novel design as the transceiver-centric HBD architecture.
We realize this design with InfiniteHBD, which has three key components as shown in Figure 2.
Design 1: Silicon Photonics based OCS transceiver (OCSTrx) (§4.1). To enable large-scale deployment, we require a low-cost, low-power transceiver with Optical Circuit Switching (OCS) support. Unlike prior high-radix switches solutions that rely on MEMS-based switching (Urata et al., 2022a; Yeow et al., 2001), we leverage advances in Silicon Photonics (SiPh), which offer a simpler structure, lower cost, and reduced power consumption—making them well-suited for commercial transceivers.
Our SiPh-based OCS transceiver (OCSTrx), shown on the left of Figure 2, provides two types of communication paths: i) Cross-lane loopback path (path 3), enabling direct GPU-to-GPU communication within the node, which can be used to construct dynamic size topologies; ii) Dual external paths (path 1&2), connecting to external nodes. All these paths utilize time-division bandwidth allocation, featuring sub-1ms switching latency. With this capability, our OCSTrx allows dynamic reallocation of full GPU bandwidth to an active external path rather than splitting bandwidth across multiple paths. This eliminates redundant link waste—for instance, activating one external path completely disables the other, ensuring efficient bandwidth utilization.
Design 2: Reconfigurable K-Hop Ring topology (§4.2). With OCSTrx  that provides reconfigurable connections at the transceiver, the next challenge is designing the topology. A naive starting point is the the full-mesh topology (Pandey, 2012) which can provide full connectivity among all nodes using OCSTrx . However, full-mesh design requires $O(N^{2})$ links, inducing prohibitive complexity and cost. To reduce costs while maintaining near-ideal fault tolerance and performance, we prune the full-mesh topology into a K-Hop Ring topology based on traffic locality and fault non-locality (Details in §4.2). Combining the reconfigurability of OCSTrx, we propose a reconfigurable K-Hop Ring topology, shown in the middle of Figure 2, which consists of two key parts:
i) Intra-node topology: dynamic GPU-granular ring construction is enabled by activating loopback paths. For example, while $N_{1}$-$N_{3}$ physically form a line topology, activating loopback paths creates a ring between $N_{1}$’s GPUs (1–4) and $N_{3}$’s GPUs (1–4). This mechanism allows for the construction of arbitrary-sized rings at any location, supporting optimal TP group sizes for different models while effectively minimizing resource fragmentation.
ii) Inter-node fault isolation: dual external paths connect to primary and secondary neighbors (e.g., 2-Hop Ring). When a node fails (e.g., $N_{2}$), its neighbor ($N_{1}$) activates the backup path ($N_{1}$-$N_{3}$) to bypass the fault while maintaining full bandwidth, approaching node-level fault explosion radius. §4.2 generalizes this design to $K>2$.
Design 3: HBD-DCN Orchestration Algorithm (§4.3). Designing an optimal HBD topology is crucial, but end-to-end training performance also depends on the efficient coordination between HBD and DCN. For instance, improper orchestration of TP groups can cause DP traffic to span across ToRs, resulting in DCN congestion. However, existing methods lack the ability to jointly optimize HBD and DCN coordination to alleviate congestion and enhance communication efficiency. To address this, we propose the HBD-DCN Orchestrator, as shown on the right side of Figure 2. The orchestrator takes three inputs: the user-defined job scale and parallelism strategy, the DCN topology and traffic pattern, and the real-time HBD fault pattern. It then generates the TP placement scheme, which maximizes GPU utilization and minimizes cross-ToR communication within the DCN.
4. InfiniteHBD Design
This section first introduces the innovative design of OCS transceivers (OCSTrx) based on Silicon Photonics (SiPh) chips (§4.1), a key enabler for InfiniteHBD, providing both cost efficiency and reconfigurability. Next, we present the DC-scale InfiniteHBD topology design (§4.2) based on OCSTrx. Finally, we outline the HBD-DCN orchestration algorithm (§4.3), designed to optimize communication efficiency for training jobs.
4.1. SiPh-based OCS transceiver (OCSTrx)
The OCSTrx is designed for reconfigurable point-to-multipoint connectivity. It incorporates a compact OCS-based switch with three Rx/Tx paths, utilizing the MZI switch  (Wikipedia, 2024) micro-structure with thermo-optic (TO) effect (Hauffe and Petermann, 2006) phase arms. This OCS-based switch is seamlessly integrated into the Photonic Integrated Chip (PIC) of the transceiver, serving as the MZI switch matrix within the Tx light path, and providing photodetector (PD) modules for each Rx paths.
SiPh-Based OCS. Currently, there are two predominant technological approaches for OCS. Micro Electromechanical systems (MEMS) (Urata et al., 2022a; Yeow et al., 2001) are attractive for commercial adoption due to supporting large port radix, up to a $320\times 320$ matrix (Calient.AI, 2024). Another option is SiPh-based OCS. Its structure is simpler and cheaper to manufacture, the limitation is its radix due to optical losses in multistage light path selector. Given that the locality of traffic and external paths number of OCSTrx is only two, SiPh-based OCS offers greater advantages.
So we choose the design of MZI micro-structure (Wikipedia, 2024) based SiPh-Based OCS. The basic mechanism of MZI switch elements is controlling the phase difference between light paths in two phase arms, then directs the output light to specific ports through interference at the output combiner. TO effect is utilized for phase arm control, for better switching latency compared to MEMS.
OCS Micro-Structure Design. As shown in Figure 3(a), the initial routing decision is made by two MZI switch elements, determining whether to direct the signal through external output 1&2, or the internal loopback path. Subsequently, an internal $N\times N$ MZI switch matrix is incorporated to facilitate the cross-lane loopback mechanism, exemplified by the blue and red paths. Notably, this design can reduces stages count and light attenuation of output 1&2, while ensures consistent light attenuation for them. The design is implemented on the Photonic Integrated Circuit (PIC) chip.
Transceiver Design. In OCSTrx, Tx electrical signal is amplified by linear driver and converted to optical signal by modulators as in Figure 3. One laser is coupled into the PIC as optical source. On the receiving end, multiple photodetectors capture the Rx optical signal from all available paths separately. The output from the activated photodetector is then amplified by a linear transimpedance amplifier (TIA). OCSTrx offers significant benefits, including high compactness, low power consumption, and cost-effective mass production.
4.2. InfiniteHBD Topology
In this section, we present the InfiniteHBD topology design (Figure 2) integrating OCSTrx that allows all GPUs within datacenter to be connected in a reconfigurable K-Hop Ring topology, while supporting dynamic ring construction and high fault tolerance.
Intra-node Topology. The intra-node topology is designed for dynamic ring construction and compiles with the OCP UBB 2.0 standard ((OAI), 2023). As shown in Figure 4, one node equipped with $R$ GPUs can support $R$ bundles of OCSTrx. Each OCSTrx bundle is connected to a pair of GPUs, with one GPU linking to the upper-half SerDes and the other to the lower-half. For one group of nodes connected as one line, the two GPU paris at each end can interconnect with the OCSTrx internal loopback path, forming a GPU-level ring. As shown in Figure 2, nodes $N_{1}$ and $N_{3}$ are connected in a line, where $OCSTrx_{1}(N_{1})$ and $OCSTrx_{2}(N_{3})$ activate the cross-lane loopback path, creating a ring between the 8 GPUs of $N_{1}$ and $N_{3}$. During ring construction, only two OCSTrx bundles per node are utilized, while the remaining OCSTrx operate in loopback mode. These idle OCSTrx can be replaced with direct connections, such as DAC links, offering a trade-off between cost and reliability. Figure 5(a,b) shows a 4-GPU node with varying numbers of OCSTrx bundles. Note that the topology design in this section utilizes a 4-GPU node as an example, it can be easily scaled for 8-GPU nodes.
Figure 4. OCSTrx connection within nodes. Each block contains multiple OCSTrx as one bundle, .e.g, $8\times 800Gbps$ OCSTrx for a 6.4Tbps GPU.
Inter-node Topology. We construct the inter-node topology by pruning the full-mesh design, based on two key observations: i) Traffic locality: TP Ring-AllReduce in HBD exhibits neighbor communication patterns, eliminating the need for distant connections; ii) Fault non-locality: node-side failures typically occur independently at the node level, meaning consecutive multi-node failures follow an exponentially decaying probability. Each node provides up to $2R$ external paths, allowing us to construct a DC-scale reconfigurable $K$-Hop Ring topology ($K\leq R$) by connecting them to nodes at ±$\pm 1,...,\pm K,K\leq R$. For AllReduce communication, only two out of the $2K$ links are activated once, with the others serving as backup links for fault isolation. For example (Figure 2), if $N_{2}$ fails, $OCSTrx_{2}(N_{1})$ and $OCSTrx_{1}(N_{3})$ can switch to backup links, maintaining connectivity between $N_{1}$ and $N_{3}$ while isolating $N_{2}$’s fault. As $K$ increases, the probability of encountering an unbypassed failure rapidly decreases, which is nearly negligible for $K=3$ (detailed analysis in Appendix §C). Thus, this architecture typically achieves a node-level explosion radius. Moreover, the K-Hop Ring can be broken into the K-Hop line topology, with the trade-off of reduced fault tolerance of $2K$ nodes at two ends.
(a) 2 bundles of OCSTrx
(b) 3 bundles of OCSTrx
Figure 5. 4-GPU node with OCSTrx.
4.3. HBD-DCN Orchestration Algorithm
InfiniteHBD is designed to work with arbitrary DCN, including Rail-Optimized (NVIDIA, 2021; Qian et al., 2024) and Fat-Tree (Al-Fares et al., 2008). This section co-optimizes communication performance for both HBD and DCN in InfiniteHBD.
Problem Statement. In InfiniteHBD, GPUs communicate without routing traffic, preventing congestion at any scale. In contrast, DCNs experience inevitable congestion, leading to performance degradation. To mitigate this, we leverage traffic locality to orchestrate nodes, minimizing cross-ToR traffic. Given a job $J$ requiring $N$ nodes from an available pool of $M$ ($M\geq N$), we must select and order $N$ nodes to satisfy two requirements: (1) nodes in the same TP group should communicate via InfiniteHBD, and (2) other parallel traffic should minimize congestion. Ideally, communication remains within the same ToR, confining congestion to switch-to-node links.
(a) Orchestration scheme 1.
(b) Orchestration scheme 2.
Figure 6. Illustration for problem statement of node orchestration.
A naive approach is sorting nodes based on deployment order in InfiniteHBD, fulfilling the first requirement but not the second. As shown in Figure 6(a), this method places ($N_{1}$, $N_{2}$) in the same TP group and ($N_{1}$, $N_{3}$) in the same DP group, forcing DP traffic across ToRs. A better scheme (Figure 6(b)) eliminates cross-ToR traffic and congestion. However, considering failures and multiple parallel dimensions complicates orchestration, necessitating an efficient method.
Our key insight is to arrange nodes in InfiniteHBD based on DCN traffic locality, prioritizing appropriate network distances over minimal ones. For example, in Figure 6(b), $N_{1}$’s InfiniteHBD neighbor is $N_{3}$, despite a 3-hop network distance in DCN. We propose a two-phase solution: (1) a deployment phase defining physical connections in DCN and InfiniteHBD, and (2) a runtime phase using an algorithm to orchestrate nodes for arbitrary-scale jobs.
Input: Topology of DCN and HBD $G$, Faulty Node Set $F$, Job Information $J$.
Output: Placement scheme that satisfies job scale and minimizes cross-ToR traffic.
Create graph $G_{deploy}=<S_{deploy},E_{deploy}>=\text{Deployment-Strategy}(G)$;
Initialize $high=n_{allconstraints}$, $low=0$, $placement=\{\}$;
while  $low\leq high$ do
       $mid=\lfloor\frac{low+high}{2}\rfloor$;
       $placement=\text{Placement-Fat-Tree}(G_{deploy},mid,F,J)$;
       if $placement$ satisfies job $J$ then
             $low=mid+1$;
      else
             $high=mid-1$;
return $placement$
Algorithm 1 Orchestration For Fat-Tree
Figure 7. Illustration of the deployment phase, showing only two backup links for simplicity.
Deployment Phase. Figure 7 shows node deployment in HBD and DCN. InfiniteHBD connects nodes at a network distance of 3 (i.e., cross-ToR). In a DCN with $r$ nodes per ToR, node $N_{n}$ connects to $N_{n\pm r}$ as main links, while backup links connect to $N_{n\pm 2r}$. For $1<n\leq r$, $N_{n}$ connects to $N_{D+n-r-1}$, where $D$ is the total node count (e.g., $N_{3}$ connects to $N_{14}$). Additionally, $N_{1}$ may link to the last node, forming a ring.
Runtime Phase. Without considering DCN topology, InfiniteHBD orchestrates nodes in three steps: (1) identifying cluster faults and modeling healthy nodes as a graph, (2) using Depth-First Search to find connected components, and (3) sequentially placing TP groups within these components. Due to InfiniteHBD’s topology, each TP group forms a ring.
For real-world DCNs, topology constraints refine step (2) and (3). In Fat-Tree networks, congestion arises when (1) a TP group spans multiple Aggregation-Switch domains, or (2) GPUs within a ToR have mismatched TP group ranks, forcing DP, CP, PP, SP traffic across ToRs. Thus, we aim to localize TP groups within the same Aggregation-Switch domain and align ranks within each ToR. Our scheduling algorithm minimizes cross-ToR traffic while meeting job scale requirements via a binary search over constraint variables. Algorithm 1 outlines the approach, with full details in Appendix §D.
5. Implementation
OCSTrx: We have successfully built a test board featuring the OCS Controller chip and a pre-release Photonic Integrated Circuit (PIC) module without the MZI switch matrix, as shown in Figure 8. The Controller Chip, measuring $4mm\times 4mm$, is manufactured using a 28nm process, while the PIC, sized at $10.5mm\times 13mm$, uses a 65nm CMOS process. The evaluation board supports 8 pairs of TX/RX SerDes at each end and has been validated for compatibility with various link layer protocols, including PCIe (32Gbps, 64Gbps) and Ethernet (56Gbps, 112Gbps). We assessed the power consumption of the peripheral circuitry using the test board. For an $8\times 112G$ configuration, the power consumption was 8.5 watts. With the addition of 3.2 watts for the MZI switch matrix, the overall consumption totals approximately 12 watts, meeting the QSFP-DD 800Gbps standard(Nowell et al., 2018).
Notably, the full-featured version of the PIC chip has successfully completed tape-out and is currently in the packaging and testing phase. It will be available for evaluation prior to final publication.
Small-scale Cluster: We constructed a small-scale cluster to evaluate the communication performance of the ring topology. Using 32 experimental GPUs equipped with inter-host HBD support (96 lanes on PCIe 4 protocol), we formed a physical ring utilizing fixed optical modules. This mini-cluster was manually reconfigured for both 32-GPU and 16-GPU ring topology. The communication latency and AllReduce performance is evaluated. For small packets, direct GPU-to-GPU links reduced latency by approximately 13% compared to the NVLink switch design. For large packets, the 16-GPU AllReduce utilized 77.11% of the ring bandwidth, with the utilization rate increasing to 77.26% for the 32-GPU configuration, showing minimal degradation with scaling. In comparison, the NVIDIA H100 8-GPU machine achieves an 81.77% utilization rate without SHARP. After deployment of OCSTrx, the size of communication group can be reconfigured within $1ms$, while maintaining maximum throughput.
Figure 8. Evaluation board for components of OCSTrx.
6. Large-Scale Simulation
We begin by outlining the experimental methodology and setup (§6.1). Next, we assess fault tolerance across different HBD architectures (§6.2), followed by end-to-end simulations to evaluate training performance under varying parallelism and GPU resource allocations (§6.3). We then examine the improvements in communication efficiency achieved by our orchestration algorithm (§6.4). Finally, we present a comparative cost and power analysis of different HBD architectures (§6.5). The simulations demonstrate that InfiniteHBD outperforms other architectures across all metrics.
6.1. Methodology and Setup
An in-house simulator dedicated for LLM training is used to evaluate InfiniteHBD comprehensively. The simulator supports end-to-end simulations of both model training performance and hardware faults, with the HBD-DCN orchestration algorithm seamlessly integrated into the system.
(a) TP-16.
(b) TP-32.
Figure 9. CDF of GPU waste ratio over 4-GPU node based on production fault trace.
GPU and network specification. The NVIDIA H100 (NVIDIA, 2024c) (989 TFLOPS, 80GiB) is used for the configuration of GPU in simulation. And the HBD bandwidth of GPU is set as $6.4Tbps$, which is the sum of 8 QSFP-DD OCSTrx. The DCN bandwidth is configured the same as NVIDIA ConnectX-7 ($400Gbps$). Since the simulation primarily focuses on HBD, the DCN is configured as a Fat-Tree topology (Al-Fares et al., 2008). Several HBD architectures are then evaluated, including:
  • Big-Switch: The ideal HBD design, featuring a large centralized switch with no forwarding latency that connects all nodes, as the theoretical upper limit of communication performance and fault resilience.
[…]
  • NVL-36, NVL-72, NVL-576 (NVIDIA, 2024b): HBDs with 36, 72, or 576 GPUs, GPU are interconnected via NVLink Switches.
  • TPUv4 (Jouppi et al., 2023): Centralized OCS capable of scheduling with a $4^{3}$ TPU cube granularity.
  • SiP-Ring (Khani et al., 2021): All nodes are connected in a series of static rings with fix sizes equal to the TP sizes.
GPU count per node. The simulation aligns with both 4-GPU node (.e.g. NVIDIA GB200 NVL-36/72/576 (NVIDIA, 2024b) and TPUv4 (Jouppi et al., 2023)) and 8-GPU node design (NVIDIA H100, AMD MI300X (AMD, 2025), Intel Gaudi3 (Corporation, 2023), and UBB 2.0 standard((OAI), 2023)).
(a) TP-16.
(b) TP-32.
Figure 10. GPU wastes ratio over the 4-GPU node with different GPU fault ratio based on fault model.
Parallelism strategy. Since InfiniteHBD is primarily designed for TP, the key variable is the TP size. TP-8, TP-16, TP-32, and TP-64 are tested to evaluate the fault resilience of various HBD architectures (§6.2). Additionally, other parallelism strategies, such as PP and DP, are used to simulate cross-ToR traffic and evaluate the orchestration algorithm (§6.4).
Fault patterns. The fault trace used in the simulation was collected from an 8-GPU node cluster with approximately 10K GPUs over a span of 348 days. On average, the ratio of faulty 8-GPU nodes is $2.66\%$, with the P99 value as $7.92\%$, more details in Appendix §A. In some simulations, fault traces generated based on this trace statistics are also derived.
6.2. HBD Fault Resilience
This section evaluates the fault resilience of various HBD architectures, focusing on GPU waste ratio, job fault-waiting time, and the maximum job scale supported by the cluster. The main text presents the key results, with more detailed results provided in Appendix §E.
GPU waste. Apart from faulty GPUs, issues such as fragmentation, topology disconnections, and bandwidth degradation can render healthy GPUs wasted. The GPU waste ratio quantifies the number of wasted GPUs under different fault scenarios. Figure 9 illustrates GPU waste ratios over production trace, while Figure 10 depicts the GPU waste ratio as node fault ratio vary.
GPU Num TP DP PP EP MFU
1024 16 16 4 1 0.4276
2048 16 16 8 1 0.4140
4096 32 16 8 1 0.3894
8192 32 16 16 1 0.3656
16384 64 16 16 1 0.3116
Table 5. Optimal parallelism strategies for maximize MFU of GPT-MoE under varying GPU numbers.
In these scenarios, InfiniteHBD ($K=3$) achieves near-zero GPU waste ratio, and outperforming all other architectures. Especially, the waste ratio for InfiniteHBD ($K=2$) remains almost identical to InfiniteHBD ($K=3$), allowing one bundle of OCSTrx to be saved for clusters with low fault rates. NVL-36 and NVL-72 typically experience an 11% waste ratio for TP sizes of 16 or larger, as $1/9$ of GPUs are reserved for redundant backups. NVL-576 has less fragmentation, benefiting from its larger size. TPUv4 performs well at low fault ratios and small TP sizes, but significantly degrades with larger TP sizes due to its coarse $4^{3}$ cube-based resource management, which amplifies the fault explosion radius. To sum up, InfiniteHBD demonstrates the strongest fault resilience among all architectures.
Maximum job supported. In fixed-size clusters, large job must pause when the available GPUs drop below the required count. Faced with same fault rate, cluster with lower GPU waste ratio can support larger job scales. Figure 11 shows the maximum job scale supported for various HBD architectures cluster with 9,792-GPU, simulated with the fault traces normalized for 4-GPU nodes. InfiniteHBD ($K=2$ or $K=3$) and NVL-576 lead in performance, and SiP-Ring exhibits declining efficiency as TP size increases.
Figure 11. Maximal job scale supported by 9,792 GPUs.
Job fault-waiting time. Large job must wait for the repairing when GPU availability falls below the required threshold. This simulations assume the average recovery time in the fault trace as a fixed repair duration. The total wasted time during 348 days is evaluated (Figure 12). For smaller TP sizes (TP-8/TP-16), NVL-36/NVL-72 exhibit the weakest resilience due to their 11% backup overhead. For larger TP sizes (TP-32/TP-64), SiP-Ring and TPUv4 perform worst.
6.3. Training Performance
This section analyzes the training performance of two representative large models, LLama 3.1-405B (Llama Team, AI @ Meta, 2024) and GPT-MoE (configuration detailed in Appendix §B), under various GPU resource configurations and parallelism strategies. The simulation results validate the practical applicability of the InfiniteHBD architecture. In simulations, we model practical TP and EP behaviors: For TP, increasing parallelism splits GEMMs into smaller, less efficient tasks, reducing hardware efficiency (NVIDIA, 2023); for EP, we practically set expert imbalance coefficient at 20%.
(a) TP-16.
(b) TP-32.
Figure 12. Job fault-waiting time over the 4-GPU node with different levels of job-scale.
[…]
GPT-MoE. The model utilizes the Mixture-of-Experts (MoE) architecture, with $EP\in\{1,2,4,8\}$ introduced in the simulation. Table 5 shows the optimal parallelism strategy and the corresponding MFU for GPT-MoE under various GPU resources. The optimal EP value is 1, suggesting that MoE can also achieve high efficiency with TP.
6.4. Communication Efficiency
(a) Sensitivity to cluster size.
(b) Impact of job-scale ratio.
(c) Sensitivity to fault ratio.
(d) Aggregate cost.
Figure 13. DCN traffic optimization analysis and aggregate normalized cost varies across different architectures under different fault ratios.
This section examines the impact of orchestration algorithms on DCN communication efficiency. Experiments were performed on a Fat-Tree architecture, like the setup in  (Gangidi et al., 2024b). As shown in Figure 13(a), the algorithm is not sensitive to cluster size. Therefore, the evaluation is based on TP-32 operations on InfiniteHBD with 8192 GPUs.
  • Baseline: A greedy algorithms, which randomly select nodes from the cluster and use the first permutation that meets the requirements.
  • Optimized: The HBD-DCN orchestration algorithm proposed in §4.3.
Figure 13(b) illustrates the impact of job-scale ratios (job size/total cluster GPUs) on cross-ToR traffic, where node fault ratio is 5%. Baseline consistently results in approximately 10% cross-ToR traffic. In contrast, the Optimized algorithm significantly outperforms the Baseline, reducing cross-ToR traffic to just 1.72% even at a 90% job-scale ratio. Figure 13(c) explores the sensitivity to node faults, with the job scale ratio fixed at 85%. The Baseline shows a linear increase of cross-ToR traffic, while the Optimized algorithm sustains near-zero cross-ToR traffic for fault ratios under 7%.
6.5. Cost and Power Analysis
To evaluate the interconnect costs of HBD architectures, we gather the cost and power information with the following methodologies:
  • For standard components (DAC cables, optical transceivers, fibers), pricing is sourced from official retailer websites (FS, 2024a; FIBERMALL.COM, 2024; NADDOD.COM, 2024) with a 60% wholesale discount validated against internal data.
  • For components with scarce public pricing information, such as Google Palomar OCS, NVIDIA NVLink Switch, 1.6 Tbps ACC cables/optical transceivers, the data is amalgamated from multiple sources (SemiAnalysis, 2024a, 2023, b) to enhance accuracy.
  • Public power consumption data is available for most components, though for NVLink Switch, multiple sources are combined to estimate a reasonable value.
The breakdown analysis of each architecture is provided in the Appendix §F. Based on this, the cost and power consumption are normalized according to GPU count and per-GPU bandwidth. As depicted in Table 6, InfiniteHBD exhibits the lowest interconnect cost per GPU per GBps. Under the $K=2$ configuration, its cost is only 62.84% of Google TPUv4 and 30.86% of the NVIDIA GB200 NVL-36/72, with minimal power consumption. This efficiency is primarily attributed to the avoidance of centralized switches. TPUv4 ranks second in interconnect cost and lowest in power consumption, achieved by reducing optical module use and per-port OCS costs. The NVL series has higher interconnect costs and power consumption due to its fully-connected topology and high-cost NVLink Switches. Notably, NVL-576 incurs the highest cost and power consumption due to its multilayer nonconvergent topology, which increases optical module expenses and requires more NVLink Switches.
Architecture Per-GPU Per-GPU Per-GBps
Cost Watts Cost Watts
TPUv4 1567.20 19.39 5.22 0.06
NVL-36 9563.20 75.95 10.63 0.08
NVL-72 9563.20 75.95 10.63 0.08
NVL-36x2 17924.00 150.33 19.92 0.17
NVL-576 30417.60 413.45 33.80 0.46
InfiniteHBD ($K=2$) 2626.80 48.10 3.28 0.06
InfiniteHBD ($K=3$) 3740.60 72.05 4.68 0.09
Table 6. Interconnect cost (\$) and power (watts).
Beyond interconnect costs, fault resilience variations also affect aggregate costs. The aggregate cost is defined as:
$Cost_{GPU}\times(N_{Wasted-GPU}+N_{Faulty-GPU})+Cost_{Interconnect}$
Simulations on a 10K-GPU cluster using the TP-32 configuration evaluate GPU availability under varying fault ratios across different architectures. The variation in aggregate cost for different HBD architectures under varying node fault ratios is illustrated in Figure 13(d). InfiniteHBD consistently exhibits the lowest aggregate cost. Furthermore, when the fault ratio is below 12.1%, the aggregate cost of InfiniteHBD ($K=2$) is less than that of InfiniteHBD ($K=3$), suggesting that ($K=2$) is the optimal design for most scenarios.
7. Discussion
AllToAll communication. Ring topology in InfiniteHBD struggles with AllToAll communication (e.g., EP), exhibiting poor performance at $O(p^{2})$, where $p$ is the group size. This can be improved by linking backup lines to nodes indexed at $n\pm 2^{i}$ instead of $n\pm i$ and Applying the Binary Exchange algorithm, and reduce time complexity to $O(p\log_{2}p)$. During the algorithm, OCSTrx need to connect to different GPUs with runtime switching, since OCSTrx switches in under $1ms$, reconfiguration can be overlapped with computation. For $K=2$ InfiniteHBD designs, performance matches the ideal Bruck algorithm (Bruck et al., 1997) when $p<8$. However, this design introduces complexities in construction, failover, and orchestration, and necessitates GPU routing capabilities. So it is not applied.
Simulation Scale. Simulations using real 10,000 GPU fault traces, as detailed in §6.2, were conducted on a cluster comprising 9,792 GPUs. The reason is that the simulation’s GPU count must be less than the total GPUs in the fault trace, and 9,792 is the largest number divisible by 576 and less than 10,000. This configuration allows the entire cluster to be divided into 17 NVL-576 units for the simulation.
Multi-dimension parallelism. InfiniteHBD is optimized for single-dimension parallelism. To support multi dimensional communication, two approaches are viable. 1) Independent Interconnects: Each OCSTrx bundle includes multiple OCSTrx units (e.g., 4 or 8), then link each of the unit to a separate inter-host topology. This isolates parallel dimensions but results in fixed bandwidth per dimension, leading to inefficiencies. 2) Time-Division multi-dimension: Main and backup lines of OCSTrx can be used to form separated inter-host topology, rapidly switching between them can support multi-dimension parallelism. However, this introduces complexity in managing multi-dimensional overlap and reduces the fault tolerance of InfiniteHBD.
Single-Job vs. Multi-Job. Existing studies explore multi-job scheduling in GPU clusters (Weng et al., 2022; Li et al., 2023b). Deploying certain small jobs, such as inference tasks, can mitigate GPU fragmentation. However, given the shortage of GPU in LLM training, any idle GPU—whether repurposed for small job or not—is undesirable. Thus, InfiniteHBD prioritizes single-job execution for simplicity.
OCS vs. EPS. OCSTrx enables multi-path selection, a feature also achievable with Electronic Packet Switching (EPS). For example, inter-host topology of InfiniteHBD can be implemented using UBB 2.0-based servers by adding external optical interfaces to switches in server. However, this would require twice optical modules and numerous high-throughput switching chips for the entire system, significantly increasing cost and power consumption compared to OCSTrx.
8. Related Work
HBD Architectures. HBDs are crucial for enabling communication intensive parallelism strategies (TP/EP) for LLM training. NVIDIA DGX SuperPOD (NVIDIA, 2024a) and GB200 NVL series (NVIDIA, 2024b) use any-to-any electrical switching, delivering high performance but suffering from high costs, scalability limitations, and fragmentation. In contrast, direct interconnect HBDs like Dojo (Tesla, 2024), TPUv3 (Jouppi et al., 2020), and SiP-Ring (Khani et al., 2021) improve scalability but have a large fault explosion radius. TPUv4 (Jouppi et al., 2023) and TPUv5p (Google, 2024) attempts a middle ground but still lacks full node-level fault isolation. InfiniteHBD introduces a novel architecture that reduces cost, improves scalability, minimizes fragmentation, enhances fault isolation, and dynamically supports TP.
AI DCN Architectures. MegaScale (Jiang et al., 2024a) and Meta’s (Gangidi et al., 2024b) AI DC use Clos-based topologies, while Rail-Optimized (NVIDIA, 2021) and Rail-Only (Wang et al., 2024) architectures optimize for LLM traffic patterns. Alibaba HPN (Qian et al., 2024) enhances fault tolerance with a dual-plane design. InfiniteHBD is compatible with all of them on LLM-training.
OCS Technologies. OCS enables dynamic topology reconfiguration in datacenters (Urata et al., 2022b; Jouppi et al., 2023; Liao et al., 2025). MEMS OCS-based switch supports high port counts (Urata et al., 2022a; Calient.AI, 2024), while silicon photonics (SiPh) achieves lower latency and cost (Hauffe and Petermann, 2006). This work proposes a SiPh-based OCS transceiver (OCSTrx), constructing an interconnect fabric without centralized switches.
Reconfigurable Networks. Traditional studies (Farrington et al., 2010; Wang et al., 2010; Chen et al., 2012; Porter et al., 2013; Ballani et al., 2020; Xia et al., 2015; Chen et al., 2017; Mellette et al., 2017, 2020; Hamedazimi et al., 2014; Amir et al., 2024) focus on generic DCN architectures without optimizing for LLM training traffic, leading to suboptimal topologies. Recent advancements like SiP-ML (Khani et al., 2021), TopoOpt (Wang et al., 2023), and mFabric (Liao et al., 2025) introduce dedicated training optimizations but still underutilize optical network reconfigurability for better fault tolerance and GPU utilization.
AI Job Schedulers. Schedulers such as  (Xiao et al., 2018; Mahajan et al., 2020; Gu et al., 2019; Jiang et al., 2020; Peng et al., 2019; Qiao et al., 2021) aim to improve GPU utilization. However, they exhibit dual limitations: their designs are premised on non-reconfigurable network, while also failing to consider job scheduling within HBD for optimizing traffic patterns in DCN. This work proposes a HBD-DCN orchestration algorithm based on reconfigurable networks to address these limitations.
9. Conclusion
In this paper, we propose InfiniteHBD, a novel HBD design that supports datacenter scale, dynamic TP group size and near-ideal fault explosion radius. InfiniteHBD is built upon a novel design of optical transceivers interaged with SiPh-based OCSTrx, a reconfigurable K-Hop Ring topology and a HBD-DCN orchestration algorithm to leverage the capabilities of the new hardware. Using real fault trace of 10K GPU cluster and the in-house simulator, we demonstrate that InfiniteHBD achieves GPU utilization close to the ideal model during faults, delivers superior cost and energy efficiency compared to existing designs, and provides effective control over cross-ToR DCN traffic. We believe InfiniteHBD provides an efficient scaling solution for HBD, which offers new insights for the infrastructure of the next generation Trillion-Parameters LLM training.
Appendix A Production Fault Trace
The production fault trace was collected from an 8-GPU node pretrain cluster with 10K GPUs over a period of 348 days. The trace includes details such as fault start time, fault end time, and the ID of the faulty node. Figure 14(a) and Figure 14(b) provide a macro-level overview of the production fault trace. On average, the ratio of faulty 8-GPU nodes at any given time is $2.66\%$, with a p99 value of $7.92\%$.
(a) Fault Node Ratio Trace.
(b) Cumulative Distribution.
Figure 14. Fault node trace in the production AI DC.
Since most of failure events are GPU faults, we normalized the trace of 8-GPU nodes to generate 4-GPU nodes trace. Assuming that the fault rates of GPUs are i.i.d. with a fault probability of $p$ for each GPU, and considering that a node is deemed faulty if any GPU within it fails, the fault rate of an 8-GPU node is calculated as:
$P_{fault}(8\text{-GPU})=1-(1-p)^{8}=2.66\%.$
From this, we derive $p=0.34\%$. The fault rate for a 4-GPU node is then:
$P_{fault}(4\text{-GPU})=1-(1-p)^{4}=1.34\%.$
The fault event of 4-GPU node is generate with Bayesian Equation, as:
[…]
$=\frac{P_{fault}(\text{8-GPU}\mid\text{4-GPU})P_{fault}(\text{4-GPU})}{P_{fault}(\text{8-GPU})}$
$=\frac{1\times 1.34\%}{2.66\%}=50.38\%$
Thus, whenever a fault occurs in an 8-GPU node in the original trace, each of the two corresponding 4-GPU nodes at the same location has a $50.38\%$ probability of fault. This method is used to convert the traces.
As node faults are i.i.d., the simulator linearly maps the fault trace to different network architectures.
Appendix B GPT-MoE Architecture
This model is a mixture-of-experts (MoE) model with the following configuration:
Model Configuration:
  • Number of Layers: 192
  • Inner Layer Dimension: 49152
  • Embedding Dimension: 12288
  • Hidden Dimension: 12288
  • Vocabulary Size: 64000
  • Number of Attention Heads: 128
  • Maximum Sequence Length: 2048
  • Number of Experts: 8
  • MoE Layer Ratio: 0.5
  • Top-K Experts: 2
Runtime Configuration:
  • Virtual Pipeline Parallelism: 3
  • Micro Batch Size: 1
  • Global Batch Size: 1536
  • Max Sequence Length: 2048
Appendix C Theoretical analysis of wasted GPU ratio for InfiniteHBD
The count of backup lines as $22K-$ 2 will significantly influence the fault tolerance of InfiniteHBD. We use the expectation of waste ratio caused by GPU failure and fragmentation problem to evaluate this design, the result is shown in Table 7.
For one single working node in the middle of line, the count of breakpoints $B$ on its two sides has the expectation as:
$E_{B}(\eta=1,middle)=2(P_{s}^{K}+P_{s}^{2K})$
Where $P_{s}$ is the fail probability of GPU node, and $\eta$ is count of nodes. The expectation of breakpoints count is:
Once the distance between one node and the tail of line is $\alpha<K$, it will connect to all nodes between itself and the last one, so there will be no breakpoints on this side, and the expectation of breakpoints count is less than nodes in the middle of line. Then, for any node in the line topology:
$E_{B}(\eta=1)\leq E_{B}(\eta=1,middle)$
When the distance between two nodes is $\beta\geq K$, the breakpoints among them can be calculated as independent. Once the distance $\beta<K$, as all nodes in this range are connected to these two nodes, there will be no breakpoints between them. So, the expectation is less than two independent nodes. Then,
$E_{B}(\eta=$ $2)<E_{B}(\eta=2,\beta\geq K)=2E(\eta=1)$
$E_{B}(\eta=$ $N_{s})\leq N_{s}E_{B}(\eta=1)$
For a LLM job which require a ring communication size (TP .etc) as $N_{t}$, InfiniteHBD will cut the whole line topology into several sub lines with the length of $N_{t}/R$. Once InfiniteHBD is cutting a new sub line from the remaining nodes in the line, all $N_{t}$ GPU will be wasted when one break point exist in the middle of this sub line required, shown in Figure 15. Then the expectation for waste GPU caused by one single break point is:
$E_{W}(B=1)=N_{t}R\cdot(1-(N_{t}/R)^{-1})=R(N_{t}-R)$
Figure 15. Break point can cause node waste compare to ideal situation.
As the influence between two break points only reduce the expectation of wasted GPUs, we can have this for $X$ break points:
$E_{W}(B=X)\leq XE_{W}(B=1)=XR(N_{t}-R)$
So the expectation of wasted GPU for a nodes cluster with $N_{s}$ GPU nodes is:
$E_{W}(\eta=N_{s})$ $\leq\sum P(B=X,\eta=N_{s})\cdot X\cdot E_{W}(B=1)$
$=E_{B}(\eta=N_{s})\cdot E_{W}(B=1)$
$\leq\lim_{P_{s}\rightarrow 0}2N_{s}\cdot R\cdot(N_{t}-R)P_{s}^{K}$
The final expectation of GPUs waste ratio is (1):
(1) $E_{WR}(\eta=N_{s})=\frac{E_{W}(\eta=N_{s})}{N_{g}}\leq 2(N_{t}-R)(P_{s})^{K}$
In our trace for a 348 days long pre-train job on 10K-GPU, the p99 failure rate for 8-GPU nodes is 7.92%. If a TP32 jobs is running on InfiniteHBD, we can get the upper bond for waste ratio expectation for various configuration in Table 7.
$K=2$ $K=3$ $K=4$
R=4 $9.22\%$ $0.37\%$ $1.52\times 10^{-4}$
R=8 $30.11\%$ $2.38\%$ $0.19\%$
Table 7. Upper bond for waste ratio expectation of GPU, where GPU failure rate is 1.03% and $N_{t}$ is 32
As shown in the table, for 4-GPU node ($R=4$) 3 bundles ($K=3$) design, the additional waste of GPU is less than 0.37%, while the waste ratio for $R=8,K=4$ is less than 0.19%. This is sufficient for production clusters.
Appendix D Orchestration For Fat-Tree
In this section, we introduce the orchestration algorithm under Fat-Tree DCN in detail.
Notations To ensure rigorous mathematical reasoning, we introduce the following notations:
  • $n$: number of nodes in the data-center.
  • $K$: OCSTrx bundle (see §4.2).
  • $S_{all}$: ordered set, represents all nodes numbered from 1 according to their physical connection order in DCN fabric. $|S_{all}|=n$.
  • $S$: ordered subset, represents nodes, $\forall u\in S,u\in S_{all}$. Adjacent elements in $S$ are also adjacent from the perspective of the InfiniteHBD topology.
  • $E$: The set of edges across $S$, should be equal to $\{(S_{i},S_{j})\mid 1\leq i<j\leq n,j-i\leq K\}$, representing the connections between nodes, including both primary and backup links, and $O(|E|)=O(K|S|)$.
  • $InfHBD=<S,E>$: the topology of InfiniteHBD as an undirected graph.
  • $F$: faulty nodes.
  • $HealthyHBD=<H,HE>$: healthy node subgraph where the set of healthy nodes $H=S-F$ and the edge set $HE=\{(u,v)\mid u\in H\text{ and }v\in H\text{ and }(u,v)\in E\}$.
  • $t$: TP size, number of GPUs in one TP Group.
  • $r$: GPU ranks per node.
  • $m=t/r$: number of nodes in a TP group.
  • $s$: job scale, number of GPUs required for the job.
  • $d$: Aggregation-Switches Domain size. Number of nodes under coverage of one group of Aggregation-Switches.
  • $n_{constrains}$: number of applied constraints in binary-search-based orchestration algorithm.
  • $p$: number of nodes under each ToR.
  • $l$: shortest sub-line length under fat-tree orchestration.
  • $n_{maxsubline}=\lfloor\frac{nd}{p}\rfloor$: max number of sub-lines.
  • $G_{deploy}=<S_{deploy},E_{deploy}>$: deployed topology. After applying the deployment strategy, the topology from the perspective of InfiniteHBD is described as follows: $S_{\text{deploy}}$ is an ordered set where adjacent elements correspond to adjacent nodes in InfiniteHBD, and $E_{\text{deploy}}$ represents the connections between nodes.
The orchestration algorithm (Algorithm 2) without considering DCN has the overall time complexity $3\cdot O(|H|+|HE|)=O(|S|+|E|)=O((K+1)|S|)=O(|S|)$.
Input: $\text{InfHBD}=\langle S,E\rangle$, $F$, $m$
Output: Placement scheme maximizing GPU utilization
Initialize $H=S-F$;
Initialize $HE=\{(u,v)\mid u\in H\text{ and }v\in H\text{ and }(u,v)\in E\}$;
Create subgraph $HealthyHBD=\langle H,HE\rangle$;
Initialize $component\_list=[]$;
Initialize $visited=\{\}$;
Initialize $placement\_scheme=\{\}$;
for  each node $s$ in $H$ do
       if  $s$ not in $visited$ then
             $component=Connected-Component-DFS(s,HealthyHBD,visited)$;
             Add $component.sortedinHBD()$ to $component\_list$;
for  each $component$ in $component\_list$ do
       while  $component.size()\geq m$ do
             Add $component.pop(m)$ to $placement\_scheme$;
return $placement\_scheme$
Algorithm 2 Orchestration-DCN-Free
Input: Node ordered set $S$, OCSTrx direction $K$, parallel factor $p$
Output: Deployment topology $G_{deploy}=<S_{deploy},E_{deploy}>$
Initialize ordered set $S_{deploy}=[]$;
Initialize $l=\lfloor\frac{|S|}{p}\rfloor$;
for $i$ in $10...p-$ 1 do
       for $j$ in $10...l-$ 1 do
             Add $i+j\cdot p$ to $S_{deploy}$;
Create $E_{deploy}=\{(S_{deploy}^{i},S_{deploy}^{j})|1\leq i\leq j\leq|S_{deploy}|,j-i\leq K\}$;
return $G_{deploy}=<S_{deploy},E_{deploy}>$
Algorithm 3 Deployment-Strategy
Fat-Tree topology is another common data center topology. A typical training strategy for this topology aims to maximize the bandwidth utilization under ToR (Top of Rack) Switches. Using Meta’s two-stage clos topology(Gangidi et al., 2024a) as a reference, it can be observed that there is an attempt to run CP under ToR.
Deployment Strategy: Assuming there are $p$ nodes under each ToR, nodes with the same index under each ToR are deployed along the same parallel sub-line, and the $p$ sub-lines are connected end-to-end, as shown in Figure 7. The training strategy involves running CP $p$ across the sub-lines and running TP within them.
Orchestration Constraints. To maximize the utilization of ToR bandwidth and minimize cross-ToR traffic, the fat-tree topology introduces two constraints:
  • $\bullet$
    Aggregation-Switches Domain Constraint: The coverage domian of a group of Aggregation Switches is limited, meaning that TP groups spanning across Aggregation Switches domains would result in cross-rail traffic, which should be avoided as much as possible.
  • $\bullet$
    TP Group Alignment Constraint: A CP Group consists of TP Groups across parallel sub-lines. To keep CP traffic within the ToR, the TP Groups must be aligned. If a node fails under one ToR, all nodes under that ToR are considered failed, expanding the failure radius by a factor of $p$.
Binary-Search-Based Orchestration Algorithm. Based on the constraints and deployment strategy, we develop a binary search orchestration algorithm (see Algorithm 5) that adjusts the number of satisfied constraints. The binary search first relaxes the TP Group alignment constraints within the Aggregation-Switches Domain and then relaxes the TP Group crossing constraints between Aggregation-Switch domains (see Algorithm 4). This process is monotonic.
The time complexity of Algorithm 2 is $O(|S|)$, and the complexity of Algorithm 4 is
$\sum_{i=1}^{n_{subline}}O(|S_{subline}|)=O(\sum_{i=1}^{n_{subline}}|S_{subline}|)=O(|S_{all}|)=O(n)$
Thus, the overall time complexity of Algorithm 5 is $O(n\log n)$.
Input: $G_{deploy}=<S_{deploy},E_{deploy}>$, $n_{constraints}$, $F$, $l$, $m$, $n_{maxsubline}$, $d$, $p$
Output: Placement scheme
Initialize $placement\_scheme=\{\}$;
Initialize $n_{align}=max(0,n_{constraints}-n_{maxsubline})$, $n_{subline}=min(n_{maxsubline},n_{constraints})$;
for $i$ in $0..n_{align}-1$ do
[…]
             $sid=i*d+j$;
             if $sid\in F$ then
                   $F\cup\{\lfloor\frac{sid-1}{p}\rfloor\cdot p+1..(\lfloor\frac{sid-1}{p}\rfloor+1)\cdot p\}$;
for $i$ in $1..n_{subline}$ do
       $S_{subline}=S_{deploy}.pop(l)$;
       $E_{subline}=\{(u,v)\mid u\in S_{subline}\text{ and }v\in S_{subline}\text{ and}(u,v)\in E_{subline}\}$;
       $F_{subline}=F\cap S_{subline}$;
       $placement\_scheme=placement\_scheme\cup\text{Orchestration-Ideal}(<S_{subline},E_{subline}>,F_{subline},m)$;
$E_{res}=\{(u,v)\mid u\in S_{deploy}\text{ and }v\in S_{deploy}\text{ and }(u,v)\in E_{deploy}\}$;
$F_{res}=F\cap S_{deploy}$;
$placement\_scheme=placement\_scheme\cup\text{Orchestration-Ideal}(<S_{deploy},E_{res}>,F_{res},m)$;
return $placement\_scheme$
Algorithm 4 Placement-Fat-Tree
Input: $S$, $r$, $p$, $F$, $t$, $s$, $d$, $K$.
Output: Placement scheme that satisfies job scale and minimizes cross-rail traffic.
Initialize $m=t/r$, $n=|S|$, $l=\lfloor\frac{d}{p}\rfloor$$n_{domain}=\lfloor\frac{n}{d}\rfloor$, $n_{maxsubline}=\lfloor\frac{nd}{p}\rfloor$;
Create graph $G_{deploy}=<S_{deploy},E_{deploy}>=\text{Deployment-Strategy}(S,K,p)$;
Initialize $high=n_{domain}+n_{maxsubline}$;
Initialize $low=0$;
Initialize $placement\_scheme=\{\}$;
while  $low\leq$ high do
       $mid=\lfloor\frac{low+high}{2}\rfloor$;
       $placement\_scheme=\text{Placement-Fat-Tree}(G_{deploy},mid,F,l,m,n_{maxsubline},d,p)$;
       if $|placement\_scheme|\cdot m\cdot r\geq s$ then
             $low=mid+1$;
      else
             $high=mid-1$;
if $|placement\_scheme|\cdot m\cdot r\geq s$ then
       return $placement\_scheme$
else
       return None
Algorithm 5 Orchestration-Fat-Tree
Appendix E Additional Simulation Results for Fault Resilience
This section presents additional simulation results related to §6.2. Figure 16 shows the variation of the GPU waste ratio over time under the production fault trace. Figure 17 presents the CDF data for the GPU waste ratio. Figure 18 illustrates the waste GPU ratio for different HBD architectures under various node failure rates, including the results for TP-8 to TP-64. Figure 19 shows the proportion of job-fault waiting time relative to total time for different job scales. All the aforementioned experiments include results for TP-8, TP-16, TP-32, and TP-64 configurations.
(a) TP-8.
(b) TP-16.
(c) TP-32.
(d) TP-64.
Figure 16. GPU waste ratio over production fault trace, 4 GPU node.
(a) TP-8.
(b) TP-16.
(c) TP-32.
(d) TP-64.
Figure 17. CDF of GPU waste ratio over production fault trace, 4 GPU node.
(a) TP-8.
(b) TP-16.
(c) TP-32.
(d) TP-64.
Figure 18. GPU wastes ratio with different GPU fault ratio, 4-GPU node.
(a) TP-8.
(b) TP-16.
(c) TP-32.
(d) TP-64.
Figure 19. Job fault-waiting duration with different levels of job-scale, 4 GPU node
Appendix F Detailed Cost and power consumption Analysis
In this section, Table 8 provides a detailed description of the quantity, cost, bandwidth, and power consumption of the interconnect components in various network architectures, including Google TPUv4 (Jouppi et al., 2023), NVIDIA GB200 NVL series (NVIDIA, 2024b), Alibaba HPN(Qian et al., 2024), and InfiniteHBD.
Component Quantity Unit Cost (\$) Unit Bandwidth (GBps) Unit Power (W)
Google TPUv4(Jouppi et al., 2023) with 4096 GPU, bandwidth 300GBps/GPU
OCS(Liu et al., 2023a) 48 80000 6400 108
DAC Cable(FS, 2024b) 5120 63.60 50 0.1
Optical Module(fibermall, 2024) 6144 360 50 12
Fiber(FS, 2024e) 6144 6.80 50 0
NVIDIA GB200 NVL-36(SemiAnalysis, 2024a) with 36 GPU, bandwidth 900GBps/GPU
NVLink Switch(SemiAnalysis, 2024b) 9 28000 3600 275
DAC Cable(FS, 2024c) 2592 35.60 25 0.1
NVIDIA GB200 NVL-72(NVIDIA, 2024b)(SemiAnalysis, 2024a) with 72 GPU, bandwidth 900GBps/GPU
NVLink Switch(SemiAnalysis, 2024b) 18 28000 3600 275
DAC Cable(FS, 2024c) 5184 35.60 25 0.1
NVIDIA GB200 NVL-36x2(SemiAnalysis, 2024a) with 72 GPU, bandwidth 900GBps/GPU
NVLink Switch(SemiAnalysis, 2024b) 36 28000 3600 275
DAC Cable(FS, 2024c) 6480 35.60 25 0.1
ACC Cable(SemiAnalysis, 2024b) 162 320 200 2.5
NVIDIA GB200 NVL-576(SemiAnalysis, 2024a) with 576 GPU, bandwidth 900GBps/GPU
NVLink Switch(SemiAnalysis, 2024b) 432 28000 3600 275
DAC Cable(FS, 2024c) 41472 35.60 25 0.1
Optical Module(MSA, 2022) 4608 850 200 25
Fiber(FS, 2024e) 4608 6.80 200 0
Alibaba HPN(Qian et al., 2024) with 16320 GPU, bandwidth 50GBps/GPU
EPS(NADDOD, 2024) 360 14960 6400 3145
DAC Cable(FS, 2024c) 32640 35.60 25 0.1
Optical Module(fibermall, 2024) 28800 360 50 12
Fiber(FS, 2024e) 14400 6.80 50 0
InfiniteHBD($K=2$) with 4 GPU, bandwidth 800GBps/GPU
DAC Cable(FS, 2024d) 4 199.60 200 0.1
OCSTrx 16 600 100 12
Fiber(FS, 2024e) 16 6.80 100 0
InfiniteHBD($K=3$) with 4 GPU, bandwidth 800GBps/GPU
DAC Cable(FS, 2024d) 2 199.60 200 0.1
OCSTrx 24 600 100 12
Fiber(FS, 2024e) 24 6.80 100 0
Table 8. Interconnect cost and power consumption of components used in different network architectures.